OpenStudy (anonymous):

Write the equation of the line. passes through 1,-5 and m=-3

OpenStudy (anonymous):

y-y1=m(x-x1)

OpenStudy (anonymous):

plugin the above formula

OpenStudy (anonymous):

y--5=-3(x-1)

OpenStudy (anonymous):

simplify further

OpenStudy (anonymous):

y--5=-3x+3

OpenStudy (michele_laino):

please you have to apply the subsequent formula: \[y-y _{0}=m(x-x _{0})\] where (x_0,y_0) is your point and m is your slope

OpenStudy (anonymous):

y=-3x+3+5

OpenStudy (anonymous):

y=-3x+8

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

thats it?

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

thanks

Nnesha (nnesha):

wait wait

Nnesha (nnesha):

so you guys are using point slope formula so  multiply - by -5 and then you have to "subtract 5" not add both sides be careful with signs bug